# Quantum Computing for Query Containmentof Conjunctive Queries

### v1

@@ 5.5. RQ3: Parameter Sensitivity for QAOA
We explore the parameter sensitivity of QAOA on the 2-cycle-to-$i$-chain family, given its generic scaling behavior. Again, the experiments are run on simulators. We vary the number of layers (0, 2, 4, and 6) and compare the constrained and the unconstrained variant.
@@
Figure 6. Solution probabilities for 2-cycle-to-$i$-chain, varying input size and QAOA parameters in simulation (AerSim). The constrained variant (right charts) is always superior.
Results. Figure 6 shows how the solution probability changes with the input size as we vary QAOA parameters. Again, gray crosses indicate a solution probability of zero. On the simulator, increasing the number of layers reliably yields higher solution probabilities. The QC-CQ problem is rarely solved beyond about 20 binary variables with zero layers; adding layers lifts the probability. The most effective change is to enable constraints, as the constrained variant outperforms the unconstrained variant.
Discussion for RQ3. The results confirm that our standard configuration (30 iterations, 2 layers, 500 shots) is reasonable. QAOA with constraints is very effective in reducing the solution space and thus increasing the solution probabilities. There are scenarios where we still find solutions for more than 40 binary variables, which is impressive given the exponential search space.
@@
Figure 7. Quantum devices: 2-cycle-to-$i$-chain on (a) a quantum annealer and (b+c) QAOA on gate-based QPUs.
@@ 5.6. RQ4: Experiments on Quantum Hardware
We explore how the effects observed with simulators transfer to current quantum hardware. The experiments are run on the three quantum hardware backends using the standard configuration.
@@
With quantum annealing (Figure 7(a)), we can solve larger problem instances than with QAOA, yet we are limited to QUBO problems. Here, the more recent generation of hardware (D-Wave Advantage2) outperforms the earlier generation (Advantage). Figures 7(b) [^4] The left plot in Figure 7(b) is missing a single data point (for 8 binary variables and 3 layers), due to a vendor-side timeout in connecting to the remote device. and 7(c) show QAOA experiments on devices by IQM and IBM. As in the simulation, the constrained variant outperforms the unconstrained variant. A key difference concerns the effect of increasing the number of layers. On the simulator, more layers always improved the solution probability, but here, this is not always the case.
Discussion for RQ4. We can successfully solve QC-CQ problems on real quantum hardware, albeit of smaller size. Our QAOA experiments on hardware confirm that the constrained variant is superior. On an idealized simulator, adding layers increases expressivity and can improve solution probability. Yet on real devices, deeper circuits can accumulate more gate and decoherence errors, which may cause a loss in fidelity.
@@ 6. Investigation of Energy Landscapes
To better understand the different scaling behavior of the two parameterized QC-CQ problems, we investigated their so-called energy landscapes. Recall that we aim to minimize a multidimensional binary function that is represented by the polynomial $p(\overline{x})$ for a query pair $(q_{1},q_{2})$. The energy landscape of $p(\overline{x})$ refers to the distribution of polynomial values across all possible input bitvectors (in this context called states), where (values of) neighboring bitvectors (i.e., they differ in one bit) are considered to be directly connected.
While the two problem families seem similar, the energy landscape of 2-chain-to-$i$-star exhibits distinguishing structural characteristics. First, it is shaped such that every local minimum is also a global minimum, whereas 2-cycle-to-$i$-chain exhibits an increasing number of local but not global minima. Therefore, the optimum is much more likely to be reached for the former, as the algorithms do not “get stuck” locally. Second, the energy landscape has a large plateau of zero-energy states that is directly connected to all negatively valued (and in particular, the optimal-valued) states. Due to direct connections, the optimal value can be found more frequently, which explains the high solution probabilities. In reverse, it implies that the simulated annealing algorithm can jump out of the optimum and “get lost” into the zero-plateau. The probability of escape decreases with increasing $i$ due to larger energy differences. Moreover, the relative size of the zero-plateau decreases with growing $i$, as the likelihood of “getting lost” in the plateau also decreases. This justifies our hypothesis that for small $i$, the simulated annealer can sometimes escape into the zero-plateau, leading to lower solution probabilities on smaller instances. This effect is specific for simulated annealing and not observed for (simulated) quantum annealing or QAOA. Appendix E provides further details.
@@ 7. Discussion
Quantum stage. Our experiments confirm that our approach is sound. We did not custom-tailor the QC-CQ problems to the low-level specifics of the quantum hardware (such as QPU topology), as often done in academic explorations. Instead, we used community-authored problem instances, which underlines our success.
@@
Postprocessing. Postprocessing ensures the absence of false positives. We did not observe high rates of false negatives, which constitute missed optimization opportunities in a database query engine using containment checks for query optimization. These are crucial prerequisites for a future database system integration.
@@ 8. Related Work
Query containment is a fundamental decision problem with numerous practical applications. For conjunctive queries under set semantics (the focus of our work), containment is known to be NP-complete (Chandra and Merlin, 1977). Under bag semantics, multiplicities matter and the problem becomes substantially more intricate. A characterization analogous to Theorem 2.5 is not known, and it remains open whether the query containment problem for conjunctive queries under bag semantics is decidable. Recent advances in database theory have established decidability for restricted fragments and undecidability for more expressive query languages, cf. (Konstantinidis and Mogavero, 2019; Marcinkowski and Ostropolski-Nalewaja, 2025).
The problem of deciding query equivalence is related, yet different. Recent work on checking the equivalence of restricted SQL queries encodes the problem for constraint solvers (Chu et al., 2018). In the case of inequivalence, Cosette (Chu et al., 2017) produces a data instance that serves as a counterexample. Similarly to the certifying witnesses produced by our approach, such counterexamples allow to validate the decision.
A recent line of research has been investigating quantum computing for cost-based query optimization (Schönberger et al., 2023a; Liu et al., 2025b; Franz et al., 2024; Liu et al., 2025b; Nayak et al., 2023; Trummer, 2025; Trummer and Koch, 2016; Fankhauser et al., 2023). Like these works, we also utilize QUBO and QAOA formulations, yet we target semantic query optimization for the first time.
There is a recent trend showing that formulations of database problems originally designed for quantum computing can perform remarkably well when executed on classical hardware. These quantum-inspired approaches (Schönberger et al., 2023b, 2025) demonstrate that reformulating classical problems can unlock new solution strategies. Inversely to applying quantum computing to database problems, there are proposals to apply database theory research to solve quantum computing problems, such as simulating circuits (Hai et al., 2025; Geerts and Hai, 2025).
Within the quantum algorithms community, there is an active line of research on quantum formulations for graph structured problems. This includes graph isomorphism (Calude et al., 2017; Mariella and Simonetto, 2023; Hua and Dinneen, 2020), which can be seen as a restricted form of graph homomorphism, as well as quantum notions of homomorphism (Mancinska and Roberson, 2016). While this work is related, the modeling assumptions differ. In quantum algorithms literature, graphs are typically represented by adjacency matrices. In contrast, we consider homomorphisms that arise from tableau queries. Thus, existing formulations would require substantial reformulation.
@@ 9. Conclusion
We present a constructive method for discovering a homomorphism that certifies query containment for conjunctive queries under set semantics. This formulation enables, for the first time, an experimental implementation on quantum hardware. A central contribution of our work is the explicit construction of this certifying homomorphism. As a result, our approach does not produce false positives, which is essential for integration into practical systems. Yet an actual integration into production systems remains future work. Quantum hardware is not yet mature enough for deployment in real-world database systems. Nevertheless, our results provide a proof-of-concept that tasks such as query containment can be encoded and executed on quantum devices. Thus, we take a first step toward quantum-powered semantic query optimization.
@@ Acknowledgements.
Luisa Gerlach’s contribution was funded by the Sponsor Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) https://gepris.dfg.de/gepris/projekt/414984028 – Project-ID Grant #414984028 – CRC 1404 FONDA. We are grateful to the D-Wave Launchpad program, as well as the IBM Quantum Credits program for supporting our experiments.
@@
Appendix A Description of the quantum algorithms and simulated annealing
In this section, we provide for the interested reader further details of the solution techniques that are mentioned in Section 2.2 and used to solve the discrete optimization problems formulated in Section 3.
@@ Appendix A Description of the quantum algorithms and simulated annealing
Here, $p\colon\{0,1\}^{n}\rightarrow\mathbb{R}$ is the objective function; it is a polynomial on $n$ binary variables $\overline{x}=(x_{1},\dotsc,x_{n})$ with real-valued coefficients. The search space is the set $C\subseteq\{0,1\}^{n}$, and the goal is to find a tuple $\overline{x}\in C$ such that the value $p(\overline{x})$ is as small as possible. This optimization problem is called unconstrained if $C=\{0,1\}^{n}$, and it is called constrained if $C\subsetneq\{0,1\}^{n}$.
@@
where $K$ is a finite set, $c_{k}$ is a real-valued coefficient and $I_{k}\subseteq\{1,\ldots,n\}$ for every $k\in K$. The degree of such a monomial $m_{k}(\overline{x})$ is $\deg m_{k}(\overline{x}):=|I_{k}|$, and the degree of $p$ is the maximum degree of its monomials. Since each variable $x_{i}$ can be assigned only the values 0 and 1, every linear term can also be expressed as a quadratic term by replacing a variable $x_{i}$ by $x_{i}^{2}$ (due to $0\cdot 0=0$ and $1\cdot 1=1$). In particular, if the degree of $p$ is $\leqslant 2$, this implies that we can assume without loss of generality that every monomial of $p$ has degree exactly 2; this setting is known as Quadratic Unconstrained Binary Optimization problems (for short: QUBOs). Therefore, such a polynomial can alternatively be represented by a real-valued upper triangular $n\times n$-matrix $Q$, representing the objective function
$p(\overline{x})\ :=\ \ \overline{x}^{T}\,Q\,\overline{x}\ \ =\ \ \sum_{i\leqslant j}\,q_{ij}\cdot x_{i}\cdot x_{j}\,$,
@@
A.1. Quantum Approximate Optimization Algorithm for gate-based quantum computers
The Quantum Approximate Optimization Algorithm (Farhi et al., 2014) was developed to find approximate solutions to combinatorial optimization problems. The algorithm has a hybrid quantum-classical structure in which quantum and classical components alternate. During the quantum part, QAOA evaluates the expected value of the objective function with respect to a parameter-dependent quantum state. The classical output of the measurements on the quantum circuit acts as a feedback which is processed in the classical part of the algorithm to heuristically compute new parameters for the next quantum iteration. The new parameters are computed with the objective of minimizing the expected value of the objective function, thereby increasing the probability of (near-)optimal results, which eventually converges.
The quantum circuit for QAOA is composed of layers: Each layer is defined by a function of the so-called mixing Hamiltonian $H_{m}$ which describes the initial state and the problem Hamiltonian
@@ A.1. Quantum Approximate Optimization Algorithm for gate-based quantum computers
that describes the objective function. For simplicity, one can imagine that for advancing layers in the circuit, the quantum state shifts from the initial state (i.e., the ground state – which is the eigenvector associated with the minimum eigenvalue – of the mixing Hamiltonian) towards the solution state (i.e., the ground state of problem Hamiltonian).
@@
and the mixer operator is implemented by parametrized single-qubit rotations in constant depth. For the polynomial objective function (4) the operator $U_{p}$ can be written as
@@
In general, if $p$ is a degree $d\leqslant n/2$ polynomial (in practice, the number of binary variables is usually much higher than the polynomial degree $d$), then the problem-specific operator is implemented with $\mathcal{O}(d\cdot n^{d})$ gates and circuit depth: The number of square-free monomials of degree at most $d$ is $\sum_{k=1}^{d}{n\choose k}$ which is in $\mathcal{O}(n^{d})$ for a fixed $d\leqslant n/2$, and implementing a degree $d$ momomial requires $\mathcal{O}(d)$ gates and circuit depth.
@@
For constrained optimization problems, the set of feasible solutions $C\subsetneq\{0,1\}^{n}$ is a strict subset of $\{0,1\}^{n}$. Strategies for incorporating such constraints into the QAOA circuit were introduced (Bärtschi and Eidenbenz, 2020). The main idea is that only the subset of feasible solutions is explored, and the reduced size of the search space improves optimization results. This is achieved by constructing a procedure that prepares a uniform superposition of all feasible states
@@
A suitable state preparation procedure and an according mixer operator that ensure that the QAOA circuit leaves the subspace of feasible solutions invariant, are provided by circuits of depth $\mathcal{O}(n_{1})$ (Zander et al., 2024).
@@
The number of layers refers to the number of steps between the initial state and the solution state, thus defining the granularity of the discretization. A higher number of layers per circuit is associated with an increasing approximation quality, but at the same time, higher circuit depth and more gates increase the difficulty of a reliable execution on noisy intermediate-scale (NISQ) quantum hardware, which is the kind of quantum hardware that is available nowadays. From the discussion in the previous paragraph it follows that for an arbitrary quadratic polynomial $p(\overline{x})$ of $n$ binary variables, the circuits for QAOA with $\ell$ layers have a favorable scaling of depth $\mathcal{O}(\ell{\cdot}n)$ for both the constrained and unconstrained version. In general, for a degree $d\leqslant n/2$ polynomial $p(\overline{x})$, the depth scales polynomially in the number of binary variables $n$ as $\mathcal{O}(\ell{\cdot}d{\cdot}n^{d})$. Therefore, they can be efficiently executed on a quantum computer. This is however not the case when a classical simulator is used to emulate the behavior of a quantum computer: handling quantum states of $n$ qubits requires storing and manipulating $2^{n}$ complex amplitudes.
@@
The explanation of the number of shots requires a short review on the connection between the quantum and the classical part of the algorithm. The quantum part of the algorithm returns an expected value of the objective function with respect to the state represented by the quantum circuit, which is then fed into the classical part. In order to estimate the expected value of the objective function without knowing the full probability distribution of possible results of the quantum circuit, the circuit is executed a number of times and its results are measured, leading to an estimation of the expected value as the mean of the measured results. The number of shots refers to the number of executions that are performed for this estimation.
@@
Currently available quantum hardware is referred to as Noisy Intermediate-Scale Quantum (NISQ) hardware. Only a small number of physical qubits (e.g., up to 156 qubits on state-of-the art IBM devices) are available, and computations are affected by limitations such as short coherence times that allow only shallow circuits to be run, and noise leading to gate and readout errors. In particular, generating highly entangled quantum states across multiple qubits – a key resource in quantum computation – remains challenging (Zander and Becker, 2025). To overcome these limitations, quantum error correction, which combines several physical qubits to build a logical qubit, will be required. To date, there have been encouraging demonstrations of this technology (goo, 2025; Putterman et al., 2025), yet significant challenges in hardware and software engineering need to be addressed for quantum computing to mature.
At this point, quantum hardware is accessed mainly through cloud-based services such as IBM Quantum Platform (IBM, 2025) or IQM Resonance (IQM, 2025). Hybrid algorithms such as QAOA consist of a feedback loop between quantum and classical computation. This translates to repeated round-trips between the local client and the remote quantum backend. For each iteration, quantum circuits are sent to the remote backend and measurement results are transmitted back to the local client. Except for latency issues, this also results in being re-queued in every iteration. The latter issue has been addressed by IBM with the “Session” execution mode that provides exclusive access to the QPU. Moving forward, a tighter integration between quantum and classical high-performance computing will enable hybrid programs to run seamlessly and efficiently.
Due to the aforementioned restrictions regarding reliability and limited access to quantum hardware, quantum algorithms and applications thereof are mostly studied using simulators that run on classical hardware. Hence, simulators are software tools that aim to emulate quantum computations. In general, the costs of simulating quantum computations on classical computers increase exponentially in the number of qubits: For a system of $n$ qubits the quantum state is represented by a $2^{n}$-dimensional state vector. E.g., for $n=40$ storing the state vector requires approximately 16 terabytes ($2^{40}\cdot 16$ bytes) of memory. In practice, if quantum states are sparse or factorized, state vector simulators such as the Qrisp simulator (Seidel et al., 2024) can enable efficient simulations of larger systems of qubits. Moreover, matrix product state simulators (e.g., provided by Qiskit Aer (Javadi-Abhari et al., 2024)) offer an efficient tool for simulating quantum computations that rely only on a restricted amount of entanglement (Vidal, 2003).

### v2

@@ 5.5. RQ3: Parameter Sensitivity for QAOA
We explore the parameter sensitivity of QAOA on the 2-cycle-to-$i$-chain family, given its generic scaling behavior. Again, the experiments are run on simulators. We vary the number of layers (0, 2, 4, and 6) and compare the unconstrained and the constrained variant.
@@
Figure 6. Solution probabilities for 2-cycle-to-$i$-chain, varying input size and QAOA parameters in simulation (AerSim). The constrained variants (shown right) are always superior.
Results. Figure 6 shows how the solution probability changes with the input size as we vary QAOA parameters. Again, gray crosses indicate a solution probability of zero. The QC-CQ problem is rarely solved beyond 20 binary variables with zero layers; adding layers lifts the probability. Enabling constraints is very effective.
Discussion. The results confirm that our standard configuration (30 iterations, 2 layers, 500 shots) is reasonable. QAOA with constraints is very effective in reducing the solution space and thus increasing the solution probabilities. There are scenarios where we still find solutions for more than 40 binary variables, which is impressive given the exponential search space.
@@
Figure 7. Quantum devices: 2-cycle-to-$i$-chain on (a) quantum annealers and (b+c) QAOA on gate-based QPUs.
@@ 5.6. RQ4: Experiments on Quantum Hardware
We explore how the effects observed with simulators transfer to current quantum hardware. The experiments are run on the four quantum hardware devices using the standard configuration.
@@
With quantum annealing (Figure 7(a)), we can solve larger problem instances than with QAOA, yet we are limited to QUBO problems. Here, the more recent generation of hardware (D-Wave Advantage2) outperforms the earlier generation (D-Wave Advantage). Figures 7(b) [^3] The left plot in Figure 7(b) is missing a single data point (for 8 binary variables and 3 layers), due to a vendor-side timeout when connecting to the remote device. and 7(c) show QAOA experiments on devices by IQM and IBM. As in the simulation, the constrained variant outperforms the unconstrained variant. A key difference concerns the effect of increasing the number of layers. Different from the simulator, more layers do not always improve the solution probability.
Discussion. We can successfully solve QC-CQ problems on real quantum hardware, albeit of smaller size. Our QAOA experiments on hardware confirm that the constrained variant is superior. On an idealized simulator, adding layers increases expressivity and can improve solution probability. Yet on real devices, deeper circuits can accumulate more gate and decoherence errors, which may cause a loss in fidelity.
@@ 6. Investigation of Energy Landscapes
To better understand the different scaling behavior of the two parameterized QC-CQ problems (cf. Figure 5), we investigated their energy landscapes (cf. Sec. 2.2). Recall that we aim to minimize the polynomial $p(\overline{x})$ that represents the given query pair $(q_{1},q_{2})$. The energy landscape of $p(\overline{x})$ refers to the distribution of polynomial values across all possible input bitvectors (in this context called states), where (values of) neighboring bitvectors (i.e., they differ in one bit) are considered to be directly connected.
While the two problem families seem similar at first, the energy landscape of 2-chain-to-$i$-star exhibits two distinguishing structural characteristics. First, it is shaped such that every local minimum is also a global minimum, whereas 2-cycle-to-$i$-chain exhibits an increasing number of local but not global minima. Therefore, the optimum is much more likely to be reached for the former, as the solvers do not “get stuck” locally. Second, the energy landscape of 2-chain-to-$i$-star has a large plateau of zero-energy states that is directly connected to all negatively valued (and in particular, the optimal-valued) states. Due to these connections, the optimal value can be found more frequently, which explains the high solution probabilities in Figure 5(b). For simulated annealing, however, the direct connections can cause the solver to reversely jump out of the optimum and then “get lost” in the zero-plateau. The probability for that jump decreases with growing $i$ due to larger energy differences. Moreover, the likelihood of “getting lost” in the zero-plateau also decreases, as the relative size of the zero-plateau decreases with growing $i$. Regarding Figure 5(b), we believe that for small $i$, the simulated annealer sometimes escapes into the zero-plateau, leading to the lower solution probabilities observed on smaller instances. This effect is specific to SA and not observed for (S)QA or QAOA. Appendix E provides further details.
@@ 7. Discussion
Quantum stage. Our experiments confirm that our approach is sound. We did not custom-tailor the QC-CQ problems to the low-level specifics of the quantum hardware (e.g., QPU topology), as often done in academic explorations. Rather, we included community-authored problem instances, which underlines our success.
@@
Postprocessing. Our problem formulation ensures the absence of false positives, and postprocessing produces verifiable witnesses for positive answers. We did not observe high rates of false negatives, which constitute missed opportunities in a database query engine using containment checks for query optimization. These are crucial prerequisites for a future database system integration.
@@ 8. Related Work
Query containment is a fundamental decision problem with numerous practical applications. For conjunctive queries under set semantics (the focus here), containment is known to be NP-complete (Chandra and Merlin, 1977). Under bag semantics, multiplicities matter and the problem becomes substantially more intricate. A characterization analogous to Theorem 2.5 is not known, and it remains open whether the query containment problem is decidable. Recent advances in database theory have established decidability for restricted fragments and undecidability for more expressive query languages, cf. (Konstantinidis and Mogavero, 2019; Marcinkowski and Ostropolski-Nalewaja, 2025).
The problem of deciding query equivalence is related. Recent work on equivalence of restricted SQL queries encodes the problem for constraint solvers (Chu et al., 2018). In the case of inequivalence, Cosette (Chu et al., 2017) produces a data instance serving as a counterexample. This validates the decision, similarly to the certifying witnesses in our approach.
A line of research has been investigating quantum computing for cost-based query optimization (Schönberger et al., 2023a; Liu et al., 2025b; Franz et al., 2024; Liu et al., 2025b; Nayak et al., 2023; Trummer, 2025; Trummer and Koch, 2016; Fankhauser et al., 2023). Like most of these works, we also utilize QUBO and QAOA formulations, yet we target semantic query optimization for the first time.
There is a recent trend showing that formulations of database problems originally designed for quantum computing can perform remarkably well when executed on classical hardware. These quantum-inspired approaches (Schönberger et al., 2023b, 2025) demonstrate that reformulating classical problems can unlock new solution strategies. Inversely to applying quantum computing to database problems, there are also proposals to apply database theory research to solve quantum computing problems, such as simulating quantum circuits (Hai et al., 2025; Geerts and Hai, 2025).
Within the quantum algorithms community, there is an active line of research on quantum formulations for graph structured problems. This includes graph isomorphism (Calude et al., 2017; Mariella and Simonetto, 2023; Hua and Dinneen, 2020), which can be seen as a restricted form of graph homomorphism, as well as quantum notions of homomorphisms (Mancinska and Roberson, 2016). While this work is related, the modeling assumptions differ. In quantum algorithms literature, graphs are typically represented by adjacency matrices, while we consider homomorphisms that arise from tableau queries. Thus, existing formulations would require substantial modification.
@@ 9. Conclusion
We present a constructive method for discovering a homomorphism that certifies query containment for conjunctive queries under set semantics. Our approach guarantees the absence of false positives and, for the first time, enables experimental evaluation on quantum hardware. Although current quantum devices are not yet suitable for real-world deployment, our results provide a proof of concept that query containment can be translated into an optimization problem and solved on quantum devices, taking a first step toward quantum-powered semantic query optimization.
@@ Acknowledgements.
Luisa Gerlach’s contribution was funded by the Sponsor Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) https://gepris.dfg.de/gepris/projekt/414984028 – Project-ID Grant #414984028 – CRC 1404 FONDA. We are grateful to D-Wave’s Leap Quantum LaunchPad program. We acknowledge the use of IBM Quantum Credits for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@
Appendix A Details omitted in Section 2.2
In this section, we provide for the interested reader more details of the solution techniques (quantum algorithms and simulated annealing) mentioned in Section 2.2 and used to solve the discrete optimization problems formulated in Section 3.
@@
Here, $p\colon\{0,1\}^{n}\rightarrow\mathbb{R}$ is the objective function; it is a polynomial on $n$ binary variables $\overline{x}=(x_{1},\dotsc,x_{n})$ with real-valued coefficients. The search space is the set $C\subseteq\{0,1\}^{n}$, and the goal is to find a tuple $\overline{x}\in C$ such that the value $p(\overline{x})$ is as small as possible. This optimization problem is called unconstrained if $C=\{0,1\}^{n}$, and constrained if $C\subsetneq\{0,1\}^{n}$.
@@
where $K$ is a finite set, $c_{k}$ is a real-valued coefficient and $I_{k}\subseteq\{1,\ldots,n\}$ for every $k\in K$. The degree of such a monomial $m_{k}(\overline{x})$ is $\deg m_{k}(\overline{x}):=|I_{k}|$, and the degree of $p$ is the maximum degree of its monomials. Since each variable $x_{i}$ can only be assigned the values 0 and 1, every linear term can also be expressed as a quadratic term by replacing a variable $x_{i}$ by $x_{i}^{2}$ (due to $0\cdot 0=0$ and $1\cdot 1=1$). In particular, if the degree of $p$ is $\leqslant 2$, this implies that we can assume without loss of generality that every monomial of $p$ has degree exactly 2; this setting is known as Quadratic Unconstrained Binary Optimization problems (for short: QUBOs). Therefore, such a polynomial can alternatively be represented by a real-valued upper triangular $n\times n$-matrix $Q$, representing the objective function
(5) $p(\overline{x})\ :=\ \ \overline{x}^{T}\,Q\,\overline{x}\ \ =\ \ \sum_{i\leqslant j}\,q_{ij}\cdot x_{i}\cdot x_{j}\,$,
@@
A.1. QAOA for gate-based quantum computers
The Quantum Approximate Optimization Algorithm (QAOA) (Farhi et al., 2014) was developed to find approximate solutions to combinatorial optimization problems. The algorithm has a hybrid quantum-classical structure in which quantum and classical components alternate. During the quantum part, QAOA evaluates the expected value of the objective function with respect to a parameter-dependent quantum state. The classical output of the measurements on the quantum circuit acts as a feedback which is processed in the classical part of the algorithm to heuristically compute new parameters for the next quantum iteration. The new parameters are computed with the objective of minimizing the expected value of the objective function, thereby increasing the probability of (near-)optimal results, which eventually converges.
The quantum circuit for QAOA is composed of layers: Each layer is defined by a function of the so-called mixing Hamiltonian $H_{m}$ that describes the initial state and the problem Hamiltonian
@@ A.1. QAOA for gate-based quantum computers
that describes the objective function. For simplicity, one can imagine that for advancing layers in the circuit, the quantum state shifts from the initial state (i.e., the ground state – which is the eigenvector associated with the minimum eigenvalue – of the mixing Hamiltonian) towards the solution state (i.e., the ground state of the problem Hamiltonian).
@@
and the mixer operator is implemented by parameterized single-qubit rotations in constant depth. For the polynomial objective function (4) the operator $U_{p}$ can be written as
@@
In general, if $p$ is a degree $d\leqslant n/2$ polynomial (in practice, the number of binary variables is usually much higher than the polynomial degree $d$), then the problem-specific operator is implemented with $\mathcal{O}(d\cdot n^{d})$ gates and circuit depth: The number of square-free monomials of degree at most $d$ is $\sum_{k=1}^{d}{n\choose k}$ which is in $\mathcal{O}(n^{d})$ for a fixed $d\leqslant n/2$, and implementing a degree $d$ monomial requires $\mathcal{O}(d)$ gates and circuit depth.
@@
For constrained optimization problems, the set of feasible solutions $C\subsetneq\{0,1\}^{n}$ is a strict subset of $\{0,1\}^{n}$. Strategies for incorporating such constraints into the QAOA circuit have been introduced (Bärtschi and Eidenbenz, 2020). The main idea is that only the subset of feasible solutions is explored and the reduced size of the search space improves the optimization results. This is achieved by constructing a procedure that prepares a uniform superposition of all feasible states
@@
A suitable state preparation procedure and an appropriate mixer operator ensure that the QAOA circuit leaves the subspace of feasible solutions invariant; they are provided by circuits of depth $\mathcal{O}(n_{1})$ (Zander et al., 2024).
@@
The number of layers refers to the number of steps between the initial state and the solution state, thus defining the granularity of the discretization. A higher number of layers per circuit is associated with an increasing approximation quality, but at the same time, higher circuit depth and more gates increase the difficulty of a reliable execution on noisy intermediate-scale (NISQ) quantum hardware, which is the kind of quantum hardware that is available nowadays. From the discussion in the previous paragraph it follows that for an arbitrary quadratic polynomial $p(\overline{x})$ of $n$ binary variables, the circuits for QAOA with $\ell$ layers have a favorable scaling of depth $\mathcal{O}(\ell{\cdot}n)$ for both the constrained and unconstrained version. In general, for a degree $d\leqslant n/2$ polynomial $p(\overline{x})$, the depth scales polynomially in the number of binary variables $n$ as $\mathcal{O}(\ell{\cdot}d{\cdot}n^{d})$. Hence, they can be efficiently executed on a quantum computer. This is, however, not the case when a classical simulator is used to emulate the behavior of a quantum computer: handling quantum states of $n$ qubits requires storing and manipulating $2^{n}$ complex amplitudes.
@@
The explanation of the number of shots requires a short review of the connection between the quantum and the classical part of the algorithm. The quantum part of the algorithm returns an expected value of the objective function with respect to the state represented by the quantum circuit, which is then fed into the classical part. In order to estimate the expected value of the objective function without knowing the full probability distribution of possible results of the quantum circuit, the circuit is executed a number of times and its results are measured, leading to an estimation of the expected value as the mean of the measured results. The number of shots refers to the number of executions that are performed for this estimation.
@@
Currently available quantum hardware is referred to as Noisy Intermediate-Scale Quantum (NISQ) hardware. Only a small number of physical qubits (e.g., up to 156 qubits on state-of-the art IBM devices) are available, and computations are affected by limitations such as short coherence times that allow only shallow circuits to be run and noise leading to gate and readout errors. In particular, generating highly entangled quantum states across multiple qubits – a key resource in quantum computation – remains challenging (Zander and Becker, 2025). To overcome these limitations, quantum error correction, which combines several physical qubits to build a logical qubit, will be required. To date, there have been encouraging demonstrations of this technology (goo, 2025; Putterman et al., 2025), yet significant challenges in hardware and software engineering need to be addressed for quantum computing to mature.
At this point, quantum hardware is accessed mainly through cloud-based services such as IBM Quantum Platform (IBM, 2025) or IQM Resonance (IQM, 2025). Hybrid algorithms such as QAOA consist of a feedback loop between quantum and classical computation. This translates to repeated round-trips between the local client and the remote quantum backend. For each iteration, quantum circuits are sent to the remote backend and measurement results are transmitted back to the local client. In addition to latency issues, this also results in being re-queued in every iteration. The latter issue has been addressed by IBM with the “Session” execution mode that provides exclusive access to the QPU. Moving forward, a tighter integration between quantum and classical high-performance computing will enable hybrid programs to run seamlessly and efficiently.
Due to the aforementioned restrictions regarding reliability and limited access to quantum hardware, quantum algorithms and applications thereof are mostly studied using simulators that run on classical hardware. Simulators are software tools that aim to emulate quantum computations. In general, the costs of simulating quantum computations on classical computers increase exponentially in the number of qubits: For a system of $n$ qubits the quantum state is represented by a $2^{n}$-dimensional state vector. E.g., for $n=40$ storing the state vector requires approximately 16 terabytes ($2^{40}\cdot 16$ bytes) of memory. In practice, if quantum states are sparse or factorized, state vector simulators such as the Qrisp simulator (Seidel et al., 2024) can enable efficient simulations of larger systems of qubits. Moreover, matrix product state simulators (e.g., provided by Qiskit Aer (Javadi-Abhari et al., 2024)) offer an efficient tool for simulating quantum computations that rely only on a restricted amount of entanglement (Vidal, 2003).